# Reinforcement Learning in Vision: A Survey

### v2

@@ 4.5 Evaluation of Vision Language Action Models
ImageReward (Xu et al., 2023) Apr 2023 Tr&Te Human-ranked pairs for T2I; reward model outputs scalar score (used in (Wu et al., 2025b; Xiao et al., 2024; Zhou et al., 2025b))
HPS (Wu et al., 2023b) Mar 2023 Tr&Te Human-ranked T2I pairs from diverse generation models (used in (Wu et al., 2023b))
HPS V2 (Wu et al., 2023a) Jun 2023 Tr&Te Human-ranked T2I pairs across diverse prompts, preference-based reward (used in (Duan et al., 2025; Gu et al., 2024))
Pick-a-Pic (Kirstain et al., 2023) May 2023 Tr&Te User preferences for pairwise rankings over T2I generations (used in (Wallace et al., 2024; Li et al., 2024b; Gu et al., 2024; Lee et al., 2024))
VideoReward (Liu et al., 2025b) Jan 2025 Te Human-ranked video pairs across quality, motion, and text alignment (used in (Liu et al., 2025b))
T2I-CompBench (Huang et al., 2023) Jul 2023 Tr&Te Compositional text-to-image dataset covering attributes and object relations. (used in (Gupta et al., 2025; Duan et al., 2025))
StarVector (Rodriguez et al., 2025a) Dec 2023 Tr&Te SVG Code Generation Data, Match Reward (used in (Rodriguez et al., 2025b))
AnimeReward (Zhu et al., 2025a) Apr 2025 Tr Multi-dimensional (e.g., character consisten) human preference anime videos (used in (Zhu et al., 2025a))
VideoPrefer (Wu et al., 2024) Dec 2024 Tr MLLM-annotated 135K video preference pairs (used in (Wu et al., 2024))
Table 5: Public benchmarks of VLAs most commonly used in visual RL. Only benchmarks relevant to visual reinforcement learning are included (RL-focused training & evaluation). ‘Tr’ and ‘Te’ refer to the ‘Train’ and ‘Test’, respectively.
GUI-R1-3K (Luo et al., 2025) Apr 2025 Tr&Te GUI trajectories spanning Windows, Linux, macOS, Android, and Web platforms
SE-GUI-3k (Yuan et al., 2025b) May 2025 Tr 3,018 examples (desktop / web / mobile) with instruction and bounding box
UI-R1 (Lu et al., 2025) May 2025 Tr 136 mobile GUI tasks covering 5 action types (click, scroll, swipe, text-input)
CAGUI (Zhang et al., 2025j) Jun 2025 Te 55 K trajectories from 30 Chinese Android apps, 8 domains
Mobile-R1 (Gu et al., 2025) Jun 2025 Tr&Te More than 500 online task trajectories from 28 Chinese apps
Mind2web (Deng et al., 2023) Jun 2023 Tr&Te 2 k tasks on 137 real websites; success/fail reward for RL
AITZ (Zhang et al., 2024b) Jun 2023 Tr&Te 18,643 Android screen–action pairs with CoAT reasoning
Omniac (Kapoor et al., 2024) Feb 2024 Tr&Te Desktop + web  9.8 k scripted tasks
GUICours (Chen et al., 2024b) Jun 2024 Tr&Te GUIEnv/Act/Chat datasets (10 M OCR + 67 k navigation)
Habitat (Puig et al., 2023) Oct 2023 Tr&Te Interactive embodied-AI scenes with humans and robots
VLN-CE (Krantz et al., 2020) Apr 2020 Tr&Te Continuous embodied navigation dataset with language instructions
RLBench (James et al., 2020) Apr 2020 Tr&Te Multi-task simulated robot manipulation benchmark dataset
RoboCasa (Nasiriany et al., 2024) Jun 2024 Tr Large-scale kitchen-task simulation for generalist robotics
LIBERO (Liu et al., 2023a) Jun 2023 Tr Lifelong robot learning benchmark with 100 manipulation tasks
VLABench (Zhang et al., 2024d) Dec 2024 Tr&Te Long-horizon language-conditioned manipulation benchmark for robots
@@ 4.6 Benchmarks
A variety of new benchmarks explicitly support RL-based training and evaluation in the visual domain (see Table 3). For MLLM, recent datasets target complex multi-hop reasoning and alignment with human preferences. For example, SEED-Bench-R1 (Chen et al., 2025c) introduces a hierarchical egocentric video question-answering benchmark with 5050k training questions and a human-verified validation set. Long Video RL (Chen et al., 2025f) scales up multi-step reasoning on long videos: it provides 5252k QA pairs with detailed reasoning annotations. Another recent benchmark, Ego-R1 Bench (Tian et al., 2025), focuses on ultra-long (week-long) egocentric videos; an RL-trained “chain-of-tool-thought” agent must invoke perception tools in  $7$ sequential steps on average to answer each query, illustrating the use of step-wise reasoning accuracy as a core challenge. In the image domain, VisuLogic (Xu et al., 2025a) contains $1,000$ carefully crafted visual reasoning puzzles (e.g.,   spatial and logic problems) to evaluate pure vision-centric reasoning, and most models perform only slightly above random on this benchmark.
@@
Benchmarks for visual generation tasks predominantly supply human preference data that serve as reward models for policy optimization. Datasets like ImageReward (Xu et al., 2023) and HPS v1 & v2 (Wu et al., 2023b; b) collect human-ranked pairs of text-to-image outputs, allowing one to train a scalar reward function that scores generations. Such reward models have been used to refine text-to-image diffusion models via RLHF, aligning outputs with human aesthetic preferences. Similarly, Pick-a-Pic (Kirstain et al., 2023) and VideoReward  (Liu et al., 2025b)extend this to broader user preferences (motion smoothness, text alignment). Some benchmarks also facilitate robust evaluation of generative RL agents on generalization. T2I-CompBench (Huang et al., 2023) is a text-to-image compositionality test set that requires correctly binding novel combinations of attributes and object relations – a measure of compositional generalization often used to assess RL-trained generators. Likewise, domain-specific benchmarks define verifiable success criteria as rewards: StarVector (Rodriguez et al., 2025a) provides SVG code-generation tasks with a strict shape-matching reward, and AnimeReward (Zhu et al., 2025a) targets consistency in animated video generations with multi-dimensional human preference scores (image-video coherence, character consistency, etc.).
@@
For vision–language action agents, numerous benchmarks provide expert trajectories and simulated environments with clear reward signals for policy training and robust evaluation. Many are centered on GUI and web interaction tasks, where success can be unambiguously measured. For instance, GUI-R1-3K (Luo et al., 2025) compiles $3,000+$ GUI manipulation trajectories across Windows, Linux, macOS, Android, and web platforms. It introduces an “R1-style” dense reward scheme mapping each correct action (e.g., clicking the right button, entering correct text, a valid scroll) to positive feedback, providing step-wise reinforcement that guides an agent through multi-step UI tasks. Building on this, SE-GUI (Yuan et al., 2025b) curates 3k high-quality GUI examples with grounded instructions and bounding-box annotations, which are used to train agents with a self-imitation RL strategy. Evaluation-focused suites like UI-R1 (Lu et al., 2025) define a fixed set of unseen tasks (e.g.,   136 mobile GUI tasks spanning click, scroll, swipe, text-input actions) to test generalization of learned policies. Meanwhile, web interaction benchmarks such as Mind2Web (Deng et al., 2023) offer  2,000 tasks on real websites with a binary success/failure reward for completing each task. Some datasets emphasize exact match and reproducibility: AITZ (Zhang et al., 2024b) (Android Interaction w/ CoAT reasoning) logs  18k screen-action pairs with corresponding tool-assisted rationales, and uses an exact action match reward to ensure precise adherence to instructions. On the other hand, broader benchmarks like OmniAct (Kapoor et al., 2024) and GUICoURS (Chen et al., 2024b) target generalist agent capabilities across diverse domains. OmniAct integrates nearly 10k scripted desktop and web tasks into a single environment, while GUICoURS combines multimodal resources (10M OCR observations, 67k navigation demonstrations) spanning GUI, web, and chat interfaces. The reward structures in these benchmarks are carefully crafted, from rule-based metrics to preference scores, to guide policy learning and reflect task goals. They enable visual RL agents to learn from meaningful feedback and be evaluated not just on task success, but also on alignment with human reasoning and performance on complex, long-horizon decisions.
@@ 5.2 Long-Horizon RL in VLA
Long-horizon vision–language agents (VLA) must execute dozens of atomic actions (e.g., clicks, drags, text edits) before any end-task reward is observed. Existing works such as OS-World (Abhyankar et al., 2025) and ARPO (Lu et al., 2025a) therefore fall back on sparse reward for each click and a binary task success flag yet empirical results suggest that even GRPO yields limited gains under such supervision. Future research should (i) discover intrinsic sub-goals: segment trajectories via state-change detection or language-conditioned clustering, then assign dense rewards to sub-goal completions; (ii) learn affordance critics: train contrastive vision–language models to score how much an action reduces the distance to the verbal goal, providing shaped feedback without manual labels; (iii) hierarchical or option-based RL: couple a high-level language planner that proposes semantic sub-tasks with a low-level policy fine-tuned by off-policy RL or decision transformers;
@@ 5.3 RL for Thinking with Vision
Recent works for visual planning, such as Chain-of-Focus (Zhang et al., 2025g) and Openthinkimg (Su et al., 2025b) all treat the picture as an external workspace: the agent may crop, sketch, highlight or insert visual tokens before emitting the next language token. While early prototypes rely on supervised heuristics for these spatial actions, moving to reinforcement learning exposes four open problems. (i) Action-space design. Cropping or doodling is naturally continuous ($x,y,w,h,\dots$) yet RL libraries and GPU memories favor small discrete sets. Hybrid schemes that learn a differentiable proposal policy and then refine coordinates via policy gradient fine-tuning, as hinted by BRPO (Chu et al., 2025) and VRAG-RL (Wang et al., 2025f), remain largely unexplored. (ii) Credit assignment. Most benchmarks only reward the final task success (e.g., answer correctness in VILASR (Wu et al., 2025a)); the whole visual chain-of-thought shares a single sparse scalar. Future work should mine step-wise proxy rewards, e.g., CLIP similarity increase after a crop, or entropy drop in a learned belief state—to enable bootstrapped or hierarchical RL. (iii) Data efficiency. Sketching or inserting patches triggers extra forward passes through the vision encoder, making naive on-policy RL prohibitively expensive. Relabeling (DeepEyes (Zheng et al., 2025)) and model-based imagination (Pixel Reasoner (Su et al., 2025a)) point to sample-efficient alternatives, but principled replay and uncertainty-aware planners for visual actions are still missing. Therefore, future directions include: learning structured visual skills (crop, zoom, draw) via skill-prior RL; devising cross-modal reward shaping that scores each edit by how much it simplifies the remaining reasoning; and curating benchmarks whose metrics expose not just final accuracy.
@@ 5.4 Reward Model Design for Visual Generation
A central obstacle for reinforcement-learning–based visual generation is the lack of a scalable and faithful reward function. Widely used handcrafted metrics such as FID (Heusel et al., 2017) offer a convenient numerical signal, yet correlate only weakly with human judgments of aesthetics, semantic fidelity, or temporal coherence, especially when the task extends beyond single-frame images. Recent learned critics, such as ImageReward (Xu et al., 2023) and HPS (Wu et al., 2023b) for images, and VideoReward (Liu et al., 2025b) for videos to bridge this gap by training on pairwise human-preference data, but each model targets a narrow modality and captures only a slice of perceptual quality (e.g., prompt alignment or visual appeal). As a result, policies optimized with PPO or GRPO often exploit loopholes in a single scalar signal, producing high-contrast artifacts, repetitive textures, or physically implausible motion that “game” the critic without improving real user satisfaction. The challenge, therefore, is to design reward models that (i) integrate complementary low-level signals (consistency, physics, geometry) with high-level human preferences, (ii) generalize across images, video and 3-D scenes, and (iii) remain robust against reward hacking while being cheap enough to update continually as user tastes evolve.

### v3

@@ 4.5 Evaluation of Vision Language Action Models
GUI-R1-3K [luo2025gui] Apr 2025 Tr&Te GUI trajectories spanning Windows, Linux, macOS, Android, and Web platforms
SE-GUI-3k [yuan2025enhancing] May 2025 Tr 3,018 examples (desktop / web / mobile) with instruction and bounding box
UI-R1 [lu2025ui] May 2025 Tr 136 mobile GUI tasks covering 5 action types (click, scroll, swipe, text-input)
CAGUI [zhang2025agentcpm] Jun 2025 Te 55 K trajectories from 30 Chinese Android apps, 8 domains
Mobile-R1 [gu2025mobile] Jun 2025 Tr&Te More than 500 online task trajectories from 28 Chinese apps
Mind2web [deng2023mind2web] Jun 2023 Tr&Te 2 k tasks on 137 real websites; success/fail reward for RL
AITZ [zhang2024android] Jun 2023 Tr&Te 18,643 Android screen–action pairs with CoAT reasoning
Omniac [kapoor2024omniact] Feb 2024 Tr&Te Desktop + web  9.8 k scripted tasks
GUICours [chen2024guicourse] Jun 2024 Tr&Te GUIEnv/Act/Chat datasets (10 M OCR + 67 k navigation)
Habitat [puig2023habitat] Oct 2023 Tr&Te Interactive embodied-AI scenes with humans and robots
VLN-CE [krantz2020beyond] Apr 2020 Tr&Te Continuous embodied navigation dataset with language instructions
RLBench [james2020rlbench] Apr 2020 Tr&Te Multi-task simulated robot manipulation benchmark dataset
RoboCasa [nasiriany2024robocasa] Jun 2024 Tr Large-scale kitchen-task simulation for generalist robotics
LIBERO [liu2023libero] Jun 2023 Tr Lifelong robot learning benchmark with 100 manipulation tasks
VLABench [zhang2024vlabench] Dec 2024 Tr&Te Long-horizon language-conditioned manipulation benchmark for robots
@@ 4.6 Benchmarks
A variety of new benchmarks explicitly support RL-based training and evaluation in the visual domain (see Table 4.5). For MLLM, recent datasets target complex multi-hop reasoning and alignment with human preferences. For example, SEED-Bench-R1 [chen2025exploring] introduces a hierarchical egocentric video question-answering benchmark with 5050k training questions and a human-verified validation set. Long Video RL [chen2025scaling] scales up multi-step reasoning on long videos: it provides 5252k QA pairs with detailed reasoning annotations. Another recent benchmark, Ego-R1 Bench [tian2025egor1chainoftoolthoughtultralongegocentric], focuses on ultra-long (week-long) egocentric videos; an RL-trained “chain-of-tool-thought” agent must invoke perception tools in  $7$ sequential steps on average to answer each query, illustrating the use of step-wise reasoning accuracy as a core challenge. In the image domain, VisuLogic [xu2025visulogic] contains $1,000$ carefully crafted visual reasoning puzzles (e.g.,   spatial and logic problems) to evaluate pure vision-centric reasoning, and most models perform only slightly above random on this benchmark.
@@
Benchmarks for visual generation tasks predominantly supply human preference data that serve as reward models for policy optimization. Datasets like ImageReward [xu2023imagereward] and HPS v1 & v2 [wu2023human, wu2023human] collect human-ranked pairs of text-to-image outputs, allowing one to train a scalar reward function that scores generations. Such reward models have been used to refine text-to-image diffusion models via RLHF, aligning outputs with human aesthetic preferences. Similarly, Pick-a-Pic [kirstain2023pick] and VideoReward  [liu2025improving]extend this to broader user preferences (motion smoothness, text alignment). Some benchmarks also facilitate robust evaluation of generative RL agents on generalization. T2I-CompBench [huang2023t2i] is a text-to-image compositionality test set that requires correctly binding novel combinations of attributes and object relations – a measure of compositional generalization often used to assess RL-trained generators. Likewise, domain-specific benchmarks define verifiable success criteria as rewards: StarVector [rodriguez2025starvector] provides SVG code-generation tasks with a strict shape-matching reward, and AnimeReward [zhu2025aligning] targets consistency in animated video generations with multi-dimensional human preference scores (image-video coherence, character consistency, etc.).
@@
For vision–language action agents, numerous benchmarks provide expert trajectories and simulated environments with clear reward signals for policy training and robust evaluation. Many are centered on GUI and web interaction tasks, where success can be unambiguously measured. For instance, GUI-R1-3K [luo2025gui] compiles $3,000+$ GUI manipulation trajectories across Windows, Linux, macOS, Android, and web platforms. It introduces an “R1-style” dense reward scheme mapping each correct action (e.g., clicking the right button, entering correct text, a valid scroll) to positive feedback, providing step-wise reinforcement that guides an agent through multi-step UI tasks. Building on this, SE-GUI [yuan2025enhancing] curates 3k high-quality GUI examples with grounded instructions and bounding-box annotations, which are used to train agents with a self-imitation RL strategy. Evaluation-focused suites like UI-R1 [lu2025ui] define a fixed set of unseen tasks (e.g.,   136 mobile GUI tasks spanning click, scroll, swipe, text-input actions) to test generalization of learned policies. Meanwhile, web interaction benchmarks such as Mind2Web [deng2023mind2web] offer  2,000 tasks on real websites with a binary success/failure reward for completing each task. Some datasets emphasize exact match and reproducibility: AITZ [zhang2024android] (Android Interaction w/ CoAT reasoning) logs  18k screen-action pairs with corresponding tool-assisted rationales, and uses an exact action match reward to ensure precise adherence to instructions. On the other hand, broader benchmarks like OmniAct [kapoor2024omniact] and GUICoURS [chen2024guicourse] target generalist agent capabilities across diverse domains. OmniAct integrates nearly 10k scripted desktop and web tasks into a single environment, while GUICoURS combines multimodal resources (10M OCR observations, 67k navigation demonstrations) spanning GUI, web, and chat interfaces. The reward structures in these benchmarks are carefully crafted, from rule-based metrics to preference scores, to guide policy learning and reflect task goals. They enable visual RL agents to learn from meaningful feedback and be evaluated not just on task success, but also on alignment with human reasoning and performance on complex, long-horizon decisions.
@@ 5.2 Long-Horizon RL in VLA
Long-horizon vision–language agents (VLA) must execute dozens of atomic actions (e.g., clicks, drags, text edits) before any end-task reward is observed. Existing works such as OS-World [abhyankar2025osworld] and ARPO [arpo] therefore fall back on sparse reward for each click and a binary task success flag yet empirical results suggest that even GRPO yields limited gains under such supervision. Future research should (i) discover intrinsic sub-goals: segment trajectories via state-change detection or language-conditioned clustering, then assign dense rewards to sub-goal completions; (ii) learn affordance critics: train contrastive vision–language models to score how much an action reduces the distance to the verbal goal, providing shaped feedback without manual labels; (iii) hierarchical or option-based RL: couple a high-level language planner that proposes semantic sub-tasks with a low-level policy fine-tuned by off-policy RL or decision transformers;
@@ 5.3 RL for Thinking with Vision
Recent works for visual planning, such as Chain-of-Focus [zhang2025chain] and Openthinkimg [su2025openthinkimg] all treat the picture as an external workspace: the agent may crop, sketch, highlight or insert visual tokens before emitting the next language token. While early prototypes rely on supervised heuristics for these spatial actions, moving to reinforcement learning exposes four open problems. (i) Action‐space design. Cropping or doodling is naturally continuous ($x,y,w,h,\dots$) yet RL libraries and GPU memories favor small discrete sets. Hybrid schemes that learn a differentiable proposal policy and then refine coordinates via policy gradient fine-tuning, as hinted by BRPO [chu2025qwen] and VRAG-RL [wang2025vrag], remain largely unexplored. (ii) Credit assignment. Most benchmarks only reward the final task success (e.g., answer correctness in VILASR [wu2025reinforcing]); the whole visual chain-of-thought shares a single sparse scalar. Future work should mine step-wise proxy rewards, e.g., CLIP similarity increase after a crop, or entropy drop in a learned belief state—to enable bootstrapped or hierarchical RL. (iii) Data efficiency. Sketching or inserting patches triggers extra forward passes through the vision encoder, making naive on-policy RL prohibitively expensive. Relabeling (DeepEyes [zheng2025deepeyes]) and model-based imagination (Pixel Reasoner [su2025pixel]) point to sample-efficient alternatives, but principled replay and uncertainty-aware planners for visual actions are still missing. Therefore, future directions include: learning structured visual skills (crop, zoom, draw) via skill-prior RL; devising cross-modal reward shaping that scores each edit by how much it simplifies the remaining reasoning; and curating benchmarks whose metrics expose not just final accuracy.
@@ 5.4 Reward Model Design for Visual Generation
A central obstacle for reinforcement-learning–based visual generation is the lack of a scalable and faithful reward function. Widely used handcrafted metrics such as FID [heusel2017gans] offer a convenient numerical signal, yet correlate only weakly with human judgments of aesthetics, semantic fidelity, or temporal coherence, especially when the task extends beyond single-frame images. Recent learned critics, such as ImageReward [xu2023imagereward] and HPS [wu2023human] for images, and VideoReward [liu2025improving] for videos to bridge this gap by training on pairwise human-preference data, but each model targets a narrow modality and captures only a slice of perceptual quality (e.g., prompt alignment or visual appeal). As a result, policies optimized with PPO or GRPO often exploit loopholes in a single scalar signal, producing high-contrast artifacts, repetitive textures, or physically implausible motion that “game” the critic without improving real user satisfaction. The challenge, therefore, is to design reward models that (i) integrate complementary low-level signals (consistency, physics, geometry) with high-level human preferences, (ii) generalize across images, video and 3-D scenes, and (iii) remain robust against reward hacking while being cheap enough to update continually as user tastes evolve.